The total charge entering a terminal is given by 1 = (10 – 10e^(-2t))mC. Calculate the current at t = 0.5s
1
Expert's answer
2021-10-27T07:01:31-0400

Given:

"q(t)=(10-10e^{-2t})\:\rm mC"

"t_1=0.5\:\rm s"


The electrical current is given by

"I(t)=\frac{dq(t)}{dt}=20e^{-2t}\:\rm mA"

At the instant "t=t_1"

"I(t_1)=20e^{-2*0.5}=7.4\:\rm mA"